please answer both parts thankyou!

1)An individual has $35,000 invested in a stock with a beta of 0.7 and another $55,000 invested in a stock with a beta of 2.5. If these are the only two investments in her portfolio, what is her portfolio's beta? Do not round intermediate calculations. Round your answer to two decimal places.
______
Answer #1

Beta of portfolio = weighted beta of stocks in portfolio.

Weight of each stock = value of stock / value of portfolio.

Weight of first stock = ($35,000 / ($35,000 + $55,000)) = 0.389.

Weight of second stock = ($55,000 / ($35,000 + $55,000)) = 0.611.

Beta of portfolio = (0.389 * 0.7) + (0.611 * 2.5)

Beta of portfolio = 1.80